OpenStudy (anonymous):

Write your answer as a power 2. Show your work and explain your steps. Computer capacity is often measured in bits and bytes. A bit is the smallest unit, which is a 1 or 0, in the computer’s memory. A byte is 2^3 bits. A megabyte (MB) is 2^20 bytes. How man bits are in a megabyte?

OpenStudy (anonymous):

2^20 * 2^3 = 2^(20+3) = 2^23